# Emergent classical gravity as stress-tensor deformed field theories

### v1

@@ V.2 Example: $T\bar{T}$-like deformation in $d$ dimensions
$\quad-(\mathfrak{p}_{j}\chi_{j}^{\mathfrak{q}_{j}}+\mathfrak{s}_{j}r_{j}^{\mathfrak{q}_{j}}+\frac{\mathfrak{q}_{j}-1}{\mathfrak{q}_{j}}\beta_{j})^{\frac{p_{j}}{2}}\Big{)}^{\frac{1}{p_{j}}},$ (34)
where $\mathfrak{p}_{j}$, $\mathfrak{q}_{j}$, and $\mathfrak{s}_{j}$ are arbitrary functions independent of $\chi$ and $r$. The EOM of the Ricci curvature tensor is
Substituting it into Eq. (34) recovers the effective deformed action in Eq. (30). As a two-dimensional example, the gravitational action is explicitly formulated as
$\quad S_{\text{grav}}^{(\lambda)}[g,\Gamma,\psi]$
By setting $p_{1}=p_{2}=1$, $\mathfrak{p}_{1}=\mathfrak{p}_{2}=\mathfrak{s}_{1}=\mathfrak{s}_{2}=l^{-1}$, $\mathfrak{q}_{1}=\mathfrak{q}_{2}=\frac{1}{2}$, and $\beta_{1}=\beta_{2}=-M^{2}/\sqrt{2}$, the action (34) can be identified with (36). By solving the EOM (21), the specific form of the Ricci curvature tensor is obtained,
$\mathcal{R}^{\mu}_{\nu}$ $=(1+\frac{2\sqrt{2}M^{2}l}{\sqrt{\chi_{+}}+\sqrt{\chi_{-}}})X^{\mu}_{\nu}+\sqrt{2}M^{2}l(\sqrt{2}M^{2}l+\sqrt{\chi_{+}}$
$\quad+\sqrt{\chi_{-}}-\frac{X}{\sqrt{\chi_{+}}+\sqrt{\chi_{-}}})\delta^{\mu}_{\nu},$ (37)
where $\chi_{\pm}=\frac{1}{2}(X\pm\sqrt{2X^{\mu}_{\nu}X^{\nu}_{\mu}-X^{2}})$. By plugging it into the gravitational action and using the definition of the flow equation (23), we have
$\partial_{\lambda}S^{(\lambda)}_{\text{EFT}}\!=\!\frac{\sqrt{2}}{2M^{2}l^{2}}\!\int\!d^{2}x\sqrt{g}\Big{[}(T_{\text{EFT}})^{\mu}_{\nu}(T_{\text{EFT}})^{\nu}_{\mu}\!-\!(T_{\text{EFT}})^{2}\Big{]},$ (38)
which corresponds to the flow equation of the $T\bar{T}$ deformation [8, 9, 10, 19].
VI Conclusion and Outlook
We present a framework where classical gravity emerges from stress-tensor deformations of quantum field theories. By rewriting the deformed partition function as a gravitational path integral, we establish an equivalence between a seed theory coupled to classical gravity and an effective field theory on a deformation-induced metric. This is different from the metric approach [17, 31], which relates $T\bar{T}$–like deformations coupled to Einstein gravity to undeformed matter coupled to Ricci-based gravity.
This framework applies to a broad class of gravitational models, including Einstein and Palatini gravity, with the deformation parameter acting as a coupling to the emergent geometry. The method is non-holographic, calculable, and provides a direct link between generic stress-energy deformations and classical gravitational action, without relying on extra dimensions or specific background geometries.
Several directions for future research remain. Quantum corrections should be incorporated to explore the validity of the deformation–gravity correspondence beyond the classical approximation. Extending the framework to include matter with spin, nontrivial topology, or finite temperature would enhance its applicability. Additionally, investigating whether similar deformations can generate other gravitational theories, such as massive or non-local gravity, would test the universality of the approach. A promising direction is to extract holographic properties from the current framework, as recently studied in [32], where gravity emerges as an effective phenomenon from QFT deformed by the stress tensor in the context of holographic renormalization group flow.
@@ Acknowledgements.
We thank Ronggen Cai, Tianhao Wu, and Lixin Xu for helpful discussions. This work was supported by NSFC Grant Nos. 12475053 and 12235016.
@@ A. Exact deformed action from perturbation method
To determine the exact form of the deformed theory, we compute the metric saddle point $g^{*}$ and express the deformed action in terms of the background metric $\hat{\gamma}$. Consider the metric perturbation $g=\hat{\gamma}+h$. By employing the techniques in [22, 23], we can expand the Einstein-Hilbert action to the second order of $h$,
$S_{\text{EH}}$ $=l^{2}\int d^{d}x\sqrt{\hat{\gamma}}\Big{[}\hat{R}-2\Lambda+\Big{(}R^{(1)}+\frac{1}{2}(\hat{R}-2\Lambda)\hat{\gamma}^{\mu\nu}h_{\mu\nu}\Big{)}$
$\quad+\Big{(}R^{(2)}+\frac{1}{2}R^{(1)}\hat{\gamma}^{\mu\nu}h_{\mu\nu}+\frac{1}{8}(\hat{R}-2\Lambda)h_{\mu\nu}(\hat{\gamma}^{\mu\nu}\hat{\gamma}^{\rho\sigma}-2\hat{\gamma}^{\mu\rho}\hat{\gamma}^{\nu\sigma})h_{\rho\sigma}\Big{)}+O(h^{3})\Big{]}.$ (1)
@@
$R^{(1)}$ $=\hat{\nabla}^{\mu}\hat{\nabla}^{\nu}h_{\mu\nu}-\hat{\Box}(\hat{\gamma}^{\mu\nu}h_{\mu\nu})-\hat{R}^{\mu\nu}h_{\mu\nu},$
$R^{(2)}$ $=\hat{R}^{\mu\rho}\hat{\gamma}^{\nu\sigma}h_{\mu\nu}h_{\rho\sigma}-\frac{1}{2}h_{\mu\nu}\hat{\gamma}^{\nu\sigma}\hat{\nabla}^{\rho}\hat{\nabla}^{\mu}h_{\rho\sigma}+\frac{1}{4}h_{\mu\nu}\hat{\gamma}^{\mu\rho}\hat{\gamma}^{\nu\sigma}\hat{\Box}h_{\rho\sigma}$
$\quad+\frac{1}{4}h_{\mu\nu}\hat{\gamma}^{\mu\nu}\hat{\gamma}^{\rho\sigma}\hat{\Box}h_{\rho\sigma}+(\text{total derivatives}).$ (2)
@@
$S_{\text{EH}}$ $=l^{2}\int d^{d}x\sqrt{\hat{\gamma}}\Big{[}\hat{R}-2\Lambda+\Big{(}\frac{1}{2}(\hat{R}-2\Lambda)\hat{\gamma}^{\mu\nu}h_{\mu\nu}-\hat{R}^{\mu\nu}h_{\mu\nu}\Big{)}$
$\quad+\Big{(}\frac{1}{4}h_{\mu\nu}(\hat{\gamma}^{\mu\rho}\hat{\gamma}^{\nu\sigma}-\hat{\gamma}^{\mu\nu}\hat{\gamma}^{\rho\sigma})\hat{\Box}h_{\rho\sigma}+\frac{1}{2}h_{\mu\nu}\hat{\gamma}^{\rho\sigma}\hat{\nabla}^{\mu}\hat{\nabla}^{\nu}h_{\rho\sigma}-\frac{1}{2}h_{\mu\nu}\hat{\gamma}^{\mu\rho}\hat{\nabla}^{\sigma}\hat{\nabla}^{\nu}h_{\rho\sigma}$
$\quad+h_{\mu\nu}[\hat{R}^{\mu\rho}\hat{\gamma}^{\nu\sigma}-\frac{1}{2}\hat{R}^{\mu\nu}\hat{\gamma}^{\rho\sigma}-\frac{1}{8}(\hat{R}-2\Lambda)(2\hat{\gamma}^{\mu\rho}\hat{\gamma}^{\nu\sigma}-\hat{\gamma}^{\mu\nu}\hat{\gamma}^{\rho\sigma})]h_{\rho\sigma}\Big{)}+O(h^{3})\Big{]}.$ (3)
@@
$\Big{[}\frac{1}{2}(\hat{\gamma}^{\mu\rho}\hat{\gamma}^{\nu\sigma}-\hat{\gamma}^{\mu\nu}\hat{\gamma}^{\rho\sigma})\hat{\Box}+\frac{1}{2}(\hat{\gamma}^{\rho\sigma}\hat{\nabla}^{\mu}\hat{\nabla}^{\nu}+\hat{\gamma}^{\mu\nu}\hat{\nabla}^{\rho}\hat{\nabla}^{\sigma})-\frac{1}{2}\hat{\gamma}^{\mu\rho}\hat{\nabla}^{\sigma}\hat{\nabla}^{\nu}-\frac{1}{2}\hat{\gamma}^{\nu\rho}\hat{\nabla}^{\sigma}\hat{\nabla}^{\mu}\Big{]}h_{\rho\sigma}$
$+\Big{[}\hat{R}^{\mu\rho}\hat{\gamma}^{\nu\sigma}+\hat{R}^{\nu\rho}\hat{\gamma}^{\mu\sigma}-\frac{1}{2}\hat{R}^{\mu\nu}\hat{\gamma}^{\rho\sigma}-\frac{1}{2}\hat{R}^{\rho\sigma}\hat{\gamma}^{\mu\nu}-\frac{1}{4}(\hat{R}-2\Lambda)(2\hat{\gamma}^{\mu\rho}\hat{\gamma}^{\nu\sigma}-\hat{\gamma}^{\mu\nu}\hat{\gamma}^{\rho\sigma})\Big{]}h_{\rho\sigma}$
$+\frac{1}{2}(\hat{R}-2\Lambda)\hat{\gamma}^{\mu\nu}-\hat{R}^{\mu\nu}-l^{-2}\lambda\hat{T}^{\mu\nu}-\tau l^{-2}\lambda h_{\rho\sigma}M^{\mu\nu\rho\sigma}=0,$ (4)
where $M^{\mu\nu\rho\sigma}=\frac{\partial\hat{T}^{\mu\nu}}{\partial\hat{\gamma}_{\rho\sigma}}+\frac{\partial\hat{T}^{\rho\sigma}}{\partial\hat{\gamma}_{\mu\nu}}+\frac{1}{2}\hat{T}^{\mu\nu}\hat{\gamma}^{\rho\sigma}+\frac{1}{2}\hat{T}^{\rho\sigma}\hat{\gamma}^{\mu\nu}$. To ensure that the higher-order terms of the metric perturbation do not affect the leading-order contribution to the stress tensor deformation and to maintain the validity of the perturbation analysis, the saddle point $h^{*}_{\mu\nu}$ should be proportional to $\lambda$ (which will be explained in detail in the next subsection). This condition requires that the background metric satisfy the vacuum Einstein field equations,
$\hat{R}^{\mu\nu}-\frac{1}{2}\hat{R}\hat{\gamma}^{\mu\nu}+\Lambda\hat{\gamma}^{\mu\nu}=0.$ (5)
@@
$S_{\text{gauge}}$ $=-\frac{l^{2}}{2}\int d^{d}x\sqrt{\hat{\gamma}}\hat{\gamma}^{\nu\rho}\hat{\nabla}^{\mu}\tilde{h}_{\mu\nu}\hat{\nabla}^{\sigma}\tilde{h}_{\rho\sigma},$
@@
$\tau\Big{[}\frac{1}{2}\hat{\Box}h^{\mu\nu}-\frac{1}{4}\hat{\gamma}^{\mu\nu}\hat{\Box}h^{\alpha}_{\alpha}+\frac{1}{2}(\hat{R}^{\mu\rho\nu\sigma}+\hat{R}^{\nu\rho\mu\sigma})h_{\rho\sigma}\Big{]}$
$\quad-l^{-2}\lambda\hat{T}^{\mu\nu}-\tau l^{-2}\lambda h_{\rho\sigma}M^{\mu\nu\rho\sigma}=0.$ (8)
The corresponding solution for auxiliary field takes the form of (9). Plugging it into the gravitational action, we obtain the leading-order contribution to the deformed action,
$S^{(\lambda)}[\hat{\gamma},\psi]$ $=\hat{S}[\hat{\gamma},\psi]+\frac{2\Lambda l^{2}\text{Vol}(\mathcal{M})}{\lambda(D-2)}$
$\quad+\frac{\lambda l^{-2}}{2}\int d^{D}xd^{D}y\sqrt{\hat{\gamma}(x)\hat{\gamma}(y)}G_{\mu\nu\rho\sigma}(x,y)\hat{T}^{\mu\nu}(x)[\hat{T}^{\rho\sigma}-\frac{1}{D-2}\hat{T}^{\alpha}_{\alpha}\hat{\gamma}^{\rho\sigma}](y)$

### v2

@@ IV.1 Example: $T\bar{T}$-like deformation in $d$ dimensions
$-(\mathfrak{p}_{j}\chi_{j}^{\mathfrak{q}_{j}}+\mathfrak{s}_{j}r_{j}^{\mathfrak{q}_{j}}+\frac{\mathfrak{q}_{j}-1}{\mathfrak{q}_{j}}\beta_{j})^{\frac{p_{j}}{2}}\Big)^{\frac{1}{p_{j}}},$ (17)
where $\{r_{j}\}$ are the eigenvalues of the Ricci tensor $\mathcal{R}^{\mu}_{\nu}$, assumed to be diagonalizable. $\mathfrak{p}_{j}$, $\mathfrak{q}_{j}$, and $\mathfrak{s}_{j}$ are arbitrary functions independent of $\chi$ and $r$. The EOMs for $\{r_{j}\}$ follow from Eq. (12a); a full derivation is provided in SM D. Substituting them into (17) recovers the effective deformed action in (14). As a significant example, the two-dimensional case of this formula is explicitly analyzed in SM D, which corresponds to the $T\bar{T}$ flow [48, 36, 13, 22].
V Nonlocal Stress–Tensor Deformations and Classical–Quantum Consistency
Beyond the classical saddle analysis, we examine the nonlocal stress–tensor deformation that underlies our framework and demonstrate its consistency at both classical and quantum levels. At the classical level, eliminating the metric perturbation $h_{\mu\nu}$ from the quadratic Einstein–Hilbert action yields a universal bilocal deformation,
$\delta_{\lambda}S\!=\!\frac{\lambda}{2l^{2}}\!\int\!d^{d}xd^{d}y\sqrt{\!g(x)g(y)}T^{\mu\nu}(x)H_{\mu\nu\rho\sigma}(x,y)T^{\rho\sigma}(y),$ (18)
where the kernel $H_{\mu\nu\rho\sigma}(x,y)=G_{\mu\nu\rho\sigma}(x,y)-1/(d-2)G_{\mu\nu\alpha\beta}(x,y)g^{\alpha\beta}(y)g_{\rho\sigma}(y)$ and the Green’s function satisfies $g^{\mu\rho}g^{\nu\sigma}\Box-R^{\mu\rho\nu\sigma}-R^{\nu\rho\mu\sigma}]_{x}G_{\rho\sigma\alpha\beta}(x,y)=\delta(x-y)\delta^{\mu}_{\alpha}\delta^{\nu}_{\beta}$. Setting $g_{\mu\nu}\to\hat{\gamma}_{\mu\nu}$ reproduces the leading nonlocal deformation (III) in linearized Einstein gravity.
To probe whether this semiclassical reorganization is compatible with quantum, we compute the leading one-loop contribution generated by the bilocal interaction (18) using the heat-kernel expansion (SM E) [33, 17, 7]. Diffeomorphism covariance forces the induced local effective action into a derivative expansion, $\delta\ln\mathcal{Z}^{(\lambda)}[g]\propto\int\sqrt{g}\,(\alpha_{0}+\alpha_{1}R+O(\nabla^{4}))$, where $\mathcal{Z}^{(\lambda)}[g]$ is the first-order deformed partition function, and the operator content is universal renormalization invariant, but the coefficients $\alpha_{0,1}$ depend on the UV regulator and renormalization scheme through local counterterms. We therefore interpret $\alpha_{0}$ and $\alpha_{1}$ as the renormalized vacuum energy and Newton coupling at a chosen reference scale, fixed by a concrete renormalization prescription (SM E).  [^8] The coefficients of local terms such as $\int\sqrt{g}$ and $\int\sqrt{g}R$ are scheme-dependent and can be shifted by local counterterms; the nonlocal kernel is fixed once the Green’s function prescription is specified. . Implementing this prescription in the proper time cutoff scheme yields
$\mathcal{Z}^{(\lambda)}_{\mathrm{scalar}}=\mathcal{Z}_{0}\Big(1-\frac{1}{16\pi G_{\text{eff}}}\int d^{4}x\sqrt{g}\,\bigl[R+O(\nabla^{4})\bigr]\Big),$ (19)
[…]
Classical–quantum consistency.
Crucially, the classical bilocal kernel (18) and the quantum effective action (19) arise from the same nonlocal stress–tensor deformation. Classically, eliminating the metric yields a bilocal interaction governing linearized gravitational response; quantum mechanically, loop corrections to this interaction generate a local gravitational action with the correct tensorial structure. The metric saddle appearing in the classical analysis, therefore, corresponds to the semiclassical (tree-level) limit of the quantum deformation. This demonstrates a coherent, logically self-consistent connection between classical geometric reorganization and quantum gravitational dynamics within a single framework of stress tensor deformation.
VI Conclusion
We have formulated a semiclassical correspondence in which stress–tensor deformations of a QFT reorganize into a gravitational action evaluated at the metric saddle. In particular, our results comprise: (i) a universal bilocal kernel equal to the inverse graviton operator reproducing linearized Einstein response; (ii) an explicit inverse construction yielding Palatini actions from specified stress–tensor deformations; and (iii) a one-loop local curvature term generated by the same deformation, linking the classical saddle to its quantum completion. The deformation is intrinsically nonlocal, governed by the inverse graviton operator and encoding massless spin-2 exchange.
This nonlocal structure lies outside the locality assumptions of the Weinberg–Witten theorem [46] and remains compatible with linearized Einstein dynamics. The nonlocal terms also generate quantum corrections to the effective action, offering a controlled setting in which aspects of quantum emergent gravity arise from stress–tensor deformations. In Lorentzian applications, the inverse operator in the kernel should be understood with a causal (linear-response) prescription, while a fully nonperturbative assessment of unitarity for the resulting nonlocal deformation lies beyond the scope of the present Letter and will be addressed elsewhere.
Several directions follow naturally. (i) Quantum gravity and holography: The induced metric flows parallel holographic renormalization [16, 35], suggesting routes toward emergent de Sitter or flat-space holography [38, 3] and offering a complementary perspective on information-theoretic aspects of gravity [32, 23]. (ii) Spacetime gluing: The consistent gluing of stress–tensor deformed QFTs suggests an analogue for gluing spacetimes [34, 26, 25], where implementing junction conditions is typically nontrivial.
@@ Acknowledgements.
We thank Ronggen Cai, Mingzhe Li, Roberto Tateo, Stefan Theisen, Tianhao Wu, and Lixin Xu for helpful discussions. We are particularly grateful to Tianhao Wu for informing us to apply the heat-kernel to enable our completely independent implementation of the main-text derivation, whereby a massive scalar field, under a deformation, yields linearized Einstein gravity. This work was supported by NSFC Grant Nos. 12475053, 12588101, and 12235016.
@@ A. Exact deformed action from perturbation method
To determine the exact form of the deformed theory, we compute the metric saddle point $g^{*}$ and express the deformed action in terms of the background metric $\hat{\gamma}$. Consider the metric perturbation $g=\hat{\gamma}+\lambda h$. By employing the techniques in [21, 2], we can expand the Einstein-Hilbert action to the second order of $h$,
$S_{\text{EH}}$ $=l^{2}\int d^{d}x\sqrt{\hat{\gamma}}\Big[\hat{R}+\Big(R^{(1)}+\frac{\lambda}{2}\hat{R}\hat{\gamma}^{\mu\nu}h_{\mu\nu}\Big)$
$\quad+\Big(R^{(2)}+\frac{\lambda}{2}R^{(1)}\hat{\gamma}^{\mu\nu}h_{\mu\nu}+\frac{\lambda}{8}\hat{R}h_{\mu\nu}(\hat{\gamma}^{\mu\nu}\hat{\gamma}^{\rho\sigma}-2\hat{\gamma}^{\mu\rho}\hat{\gamma}^{\nu\sigma})h_{\rho\sigma}\Big)+O(\lambda^{3}h^{3})\Big].$ (1)
@@
$R^{(1)}$ $=\lambda\hat{\nabla}^{\mu}\hat{\nabla}^{\nu}h_{\mu\nu}-\lambda\hat{\Box}(\hat{\gamma}^{\mu\nu}h_{\mu\nu})-\lambda\hat{R}^{\mu\nu}h_{\mu\nu},$
$R^{(2)}$ $=\lambda^{2}\hat{R}^{\mu\rho}\hat{\gamma}^{\nu\sigma}h_{\mu\nu}h_{\rho\sigma}-\frac{\lambda^{2}}{2}h_{\mu\nu}\hat{\gamma}^{\nu\sigma}\hat{\nabla}^{\rho}\hat{\nabla}^{\mu}h_{\rho\sigma}+\frac{\lambda^{2}}{4}h_{\mu\nu}\hat{\gamma}^{\mu\rho}\hat{\gamma}^{\nu\sigma}\hat{\Box}h_{\rho\sigma}$
$\quad+\frac{\lambda^{2}}{4}h_{\mu\nu}\hat{\gamma}^{\mu\nu}\hat{\gamma}^{\rho\sigma}\hat{\Box}h_{\rho\sigma}+(\text{total derivatives}).$ (2)
@@
$S_{\text{EH}}$ $=l^{2}\int d^{d}x\sqrt{\hat{\gamma}}\Big[\hat{R}+\lambda\Big(\frac{1}{2}\hat{R}\hat{\gamma}^{\mu\nu}h_{\mu\nu}-\hat{R}^{\mu\nu}h_{\mu\nu}\Big)$
$\quad+\Big(\frac{\lambda^{2}}{4}h_{\mu\nu}(\hat{\gamma}^{\mu\rho}\hat{\gamma}^{\nu\sigma}-\hat{\gamma}^{\mu\nu}\hat{\gamma}^{\rho\sigma})\hat{\Box}h_{\rho\sigma}+\frac{\lambda^{2}}{2}h_{\mu\nu}\hat{\gamma}^{\rho\sigma}\hat{\nabla}^{\mu}\hat{\nabla}^{\nu}h_{\rho\sigma}-\frac{\lambda^{2}}{2}h_{\mu\nu}\hat{\gamma}^{\mu\rho}\hat{\nabla}^{\sigma}\hat{\nabla}^{\nu}h_{\rho\sigma}$
$\quad+\lambda^{2}h_{\mu\nu}[\hat{R}^{\mu\rho}\hat{\gamma}^{\nu\sigma}-\frac{1}{2}\hat{R}^{\mu\nu}\hat{\gamma}^{\rho\sigma}-\frac{1}{8}\hat{R}(2\hat{\gamma}^{\mu\rho}\hat{\gamma}^{\nu\sigma}-\hat{\gamma}^{\mu\nu}\hat{\gamma}^{\rho\sigma})]h_{\rho\sigma}\Big)+O(\lambda^{3}h^{3})\Big].$ (3)
@@
$\lambda\Big[\frac{1}{2}(\hat{\gamma}^{\mu\rho}\hat{\gamma}^{\nu\sigma}-\hat{\gamma}^{\mu\nu}\hat{\gamma}^{\rho\sigma})\hat{\Box}+\frac{1}{2}(\hat{\gamma}^{\rho\sigma}\hat{\nabla}^{\mu}\hat{\nabla}^{\nu}+\hat{\gamma}^{\mu\nu}\hat{\nabla}^{\rho}\hat{\nabla}^{\sigma})-\frac{1}{2}\hat{\gamma}^{\mu\rho}\hat{\nabla}^{\sigma}\hat{\nabla}^{\nu}-\frac{1}{2}\hat{\gamma}^{\nu\rho}\hat{\nabla}^{\sigma}\hat{\nabla}^{\mu}\Big]h_{\rho\sigma}$
$+\lambda\Big[\hat{R}^{\mu\rho}\hat{\gamma}^{\nu\sigma}+\hat{R}^{\nu\rho}\hat{\gamma}^{\mu\sigma}-\frac{1}{2}\hat{R}^{\mu\nu}\hat{\gamma}^{\rho\sigma}-\frac{1}{2}\hat{R}^{\rho\sigma}\hat{\gamma}^{\mu\nu}-\frac{1}{4}\hat{R}(2\hat{\gamma}^{\mu\rho}\hat{\gamma}^{\nu\sigma}-\hat{\gamma}^{\mu\nu}\hat{\gamma}^{\rho\sigma})\Big]h_{\rho\sigma}$
$+\frac{1}{2}\hat{R}\hat{\gamma}^{\mu\nu}-\hat{R}^{\mu\nu}-l^{-2}\lambda\hat{T}^{\mu\nu}-l^{-2}\lambda^{2}h_{\rho\sigma}M^{\mu\nu\rho\sigma}=0,$ (4)
where $M^{\mu\nu\rho\sigma}=\frac{\partial\hat{T}^{\mu\nu}}{\partial\hat{\gamma}_{\rho\sigma}}+\frac{\partial\hat{T}^{\rho\sigma}}{\partial\hat{\gamma}_{\mu\nu}}+\frac{1}{2}\hat{T}^{\mu\nu}\hat{\gamma}^{\rho\sigma}+\frac{1}{2}\hat{T}^{\rho\sigma}\hat{\gamma}^{\mu\nu}$. To ensure that the higher-order terms of the metric perturbation do not affect the leading-order contribution to the stress tensor deformation and to maintain the validity of the perturbation analysis, the saddle point $h^{*}_{\mu\nu}$ should be proportional to $O(\lambda^{0})$ (which will be explained in detail in the next subsection). This condition requires that the background metric satisfy the vacuum Einstein field equations,
$\hat{R}^{\mu\nu}-\frac{1}{2}\hat{R}\hat{\gamma}^{\mu\nu}=0.$ (5)
@@
$S_{\text{gauge}}$ $=-\frac{\lambda^{2}l^{2}}{2}\int d^{d}x\sqrt{\hat{\gamma}}\hat{\gamma}^{\nu\rho}\hat{\nabla}^{\mu}\tilde{h}_{\mu\nu}\hat{\nabla}^{\sigma}\tilde{h}_{\rho\sigma},$
@@
$\frac{1}{2}\hat{\Box}h^{\mu\nu}-\frac{1}{4}\hat{\gamma}^{\mu\nu}\hat{\Box}h^{\alpha}_{\alpha}+\frac{1}{2}(\hat{R}^{\mu\rho\nu\sigma}+\hat{R}^{\nu\rho\mu\sigma})h_{\rho\sigma}-l^{-2}\hat{T}^{\mu\nu}-l^{-2}\lambda h_{\rho\sigma}M^{\mu\nu\rho\sigma}=0.$ (8)
The corresponding solution for auxiliary field takes the form
$h^{*}_{\mu\nu}(x)$ $=-\frac{2}{l^{2}}\int d^{d}y\sqrt{\hat{\gamma}(y)}G_{\mu\nu\rho\sigma}(x,y)\Big(\hat{T}^{\rho\sigma}(y)-\frac{1}{d-2}\hat{T}(y)\hat{\gamma}^{\rho\sigma}(y)\Big)+O(\lambda),$ (9)
Plugging it into the gravitational action, we obtain the leading-order contribution to the deformed action,
$S^{(\lambda)}[\hat{\gamma},\psi]$ $=\hat{S}[\hat{\gamma},\psi]+\frac{\lambda l^{-2}}{2}\int d^{d}xd^{d}y\sqrt{\hat{\gamma}(x)\hat{\gamma}(y)}G_{\mu\nu\rho\sigma}(x,y)\hat{T}^{\mu\nu}(x)[\hat{T}^{\rho\sigma}-\frac{1}{d-2}\hat{T}^{\alpha}_{\alpha}\hat{\gamma}^{\rho\sigma}](y)$